### v1

# Word length predicts word order: “Min-max”-ing drives language evolution

Abstract
Current theories of language propose an innate (Baker 2001; Chomsky 1981) or a functional (Greenberg 1963; Dryer 2007; Hawkins 2014) origin for the surface structures (i.e. word order) that we observe in languages of the world, while evolutionary modeling (Dunn et al. 2011) suggests that descent is the primary factor influencing such patterns. Although there are hypotheses for word order change from both innate and usage-based perspectives for specific languages and families, there are key disagreements between the two major proposals for mechanisms that drive the evolution of language more broadly (Wasow 2002; Levy 2008). This paper proposes a universal underlying mechanism for word order change based on a large tagged parallel dataset of over 1,500 languages representing 133 language families and 111 isolates. Results indicate that word class length is significantly correlated with word order crosslinguistically, but not in a straightforward manner, partially supporting opposing theories of processing, while at the same time predicting historical word order change in two different phylogenetic lines and explaining more variance than descent or language area in regression models. Such findings suggest an integrated “Min-Max” theory of language evolution driven by competing pressures of processing and information structure, aligning with recent efficiency-oriented (Levshina 2023) and information-theoretic proposals (Zaslavsky 2020; Tucker et al. 2025).
1 Introduction
Theories of language should describe not only the current state of languages, but also explain how they have developed (their history) and predict how they will change (their future). To date, however, existing theories have primarily focused on describing the present state of language, proposing that the current diversity in surface structures or patterns that we observe in languages of the world, such as for word order, have either an innate (Baker 2001; Chomsky 1981) or functional (Greenberg 1963; Dryer 2007; Hawkins 2014) source. More recent diachronic work has tried to answer the second question on how languages develop. From an “innateness” perspective, surface word orders are seen to be derived from an underlying (universally shared) structure, whereby parameters for surface realizations change over time (Carnie & Guilfoyle 2000; Crisma & Longobardi 2009; Martins & Cardoso 2018), while from a “functional” perspective surface word orders are the result of cumulative changes in linked structures (sometimes called ‘constructions’) that compete for adoption in individual languages (see Gerritsen & Stein 1992; Leino & Kuningas 2006; Barðdal & Gildea 2016).
Importantly, none of these perspectives are particularly explanatory of drivers in language evolution that affect all languages similarly. The closest observations in this regard come from evolutionary modeling, where descent from a common ancestor was found to be the most explanatory factor in predicting word order (Dunn et al. 2011; though see Jäger & Wahle 2021; Naranjo & Becker 2022 for alternate perspectives), and language processing, where word length (Piantadosi et al. 2011; Wasow 2022) and surprisal (Levy 2008) have been shown to impact the realization of word order. There are also more general claims that “efficiency” (variously defined) explains grammatical (Hawkins 2014; Levshina 2023) or semantic (Zaslavsky 2020) change. However, the fact that languages tend to display word orders like their family members does not explain the variation in word orders that we find within families, and perspectives on processing make very different predictions: the literature on word length predicts that short/light words will occur first in a sentence, while the literature on surprisal predicts that long/heavy words will occur first. Further, the literature on efficiency does not make actual testable cross-linguistic predictions about basic word order.
There are also other issues to consider, based on observation of the languages for which we have data. Typological surveys (Dryer & Haspelmath 2013; Skirgård et al. 2023; Bickel et al. 2023) reveal that over 80% of the sampled languages have a Subject-initial basic word order, while only about 10% have a Verb-initial order. Further, there are various correlations to be found between the orders of clausal elements, such that basic word order is related to how other clause types are ordered. While all languages allow variation in word order to different degrees, the word order most often found in declarative sentences can predict (to some degree) the ordering of other grammatical elements in a language (Dryer 1992, 2011).
It is not completely clear why one basic constituent order correlates better with one particular order of (e.g.) noun phrase elements than another, nor why some basic word orders are less well-represented crosslinguistically. This raises the question whether there are underlying processes that drive these observed tendencies. Some explanation clearly comes from descent, i.e. that languages tend to resemble their ancestral forms. However, the word order changes that languages undergo over time suggests other pressures in terms of how complex linguistic elements are and how these constituents are processed in the brain (Maurits & Griffiths 2014; Schouwstra & de Swart 2014, see also Piantadosi et al. 2011; Bickel et al. 2015; Ding et al. 2017; Fernandez 2016; Jaeger & Tily 2011; Tily 2010; Sinnemäki 2014).
Answers to these questions have been elusive in part due to a lack of data that would facilitate crosslinguistic investigations. This is because such investigations require data that is annotated for various word types or dependencies, which is unavailable for the majority of the world’s languages. The largest sample of languages used to investigate these questions looked at annotated corpora in 80 languages (Hahn & Xu 2022), the majority of which are from the Indo-European language family. Such data risks biasing findings toward a particular group of languages (Blasi et al. 2022), in much the same way that other scientific findings have historically been biased toward WEIRD demographics (Henrich et al. 2010), leading to recent calls for more expansive datasets to examine linguistic questions (Benítez-Burraco 2025; Levshina et al. 2023).
The current study addresses this by developing a large corpus of texts tagged for word class. With data sourced from the Parallel Bible Corpus (PBC; Mayer & Cysouw 2014), machine learning techniques allow for automatic tagging [^1] Automated part of speech tagging was validated via comparison with existing taggers for high-resource languages (SpaCy, 16 languages - avg nouns: 0.83, verbs: 0.74; Trankit, 31 languages - avg nouns: 0.83, verbs: 0.76) and gold-standard tagged data in the Universal Dependencies Treebanks (UDT; Zeman et al. 2024; 106 languages; https://universaldependencies.org). See Ring (2025) for more details. of word classes in 1,597 languages (roughly 20% of the world’s languages) that represent 133 families and 111 isolates.222‘Isolates’ being defined here as languages in the corpus that share no known family membership with other languages in the corpus. Individual corpora in the majority of these languages (over 1,500) contain more than 1,800 parallel sentences, and the smallest corpus contains more than 700 sentences. [^3] For clarity, the UDT has 160+ hand-annotated languages available for crosslinguistic investigation, but these are not parallel texts and the number of sentences in each language is variable, with some having fewer than 200 sentences in the v2.14 release. The ODIN project (and its re-packaged form via LECS: https://huggingface.co/datasets/lecslab/glosslm-corpus) has interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences. Additional cross-linguistic investigations using tagged corpora include Jing et al. (2021: 71 languages) and Hahn et al. (2020: 51 languages). This dwarfs both the linguistic data and the language diversity available from other sources and means that crosslinguistic investigations using this dataset can be confident that they are comparing similar contexts. The dataset (the tagged PBC) and explanations regarding its development, along with Python code for replicating the results in this paper, can be found at the linked repository [^4] https://github.com/lingdoc/taggedPBC and is described at greater length in Ring (2025).
This dataset can be used to investigate the crosslinguistic questions highlighted above regarding whether there is some pressure (beyond descent) that explains the evolution of word order in languages of the world. If there is indeed a mechanism that underlies the observed surface structures in present-day languages, this should also be one that reveals historical processes and allows us to predict future developments. Specifically, by deriving a novel corpus-based measure of word order (the “N1 ratio”) [^5] The N1 ratio was validated via comparison with existing classifications of languages in the tagged PBC and three typological databases: the World Atlas of Language Structures (WALS; Dryer & Haspelmath 2013; https://wals.info/), the AUTOTYP Database (Bickel et al. 2023; https://github.com/autotyp/autotyp-data), and Grambank (Skirgård et al. 2023; https://grambank.clld.org/) where the N1 ratio significantly differentiated (p ¡ 0.001) between SV/VS/free classifications in all three databases. See Ring (2025) for more details. from this dataset, the current paper investigates the relationship between basic word order and the lengths of particular word classes (nouns/arguments and verbs/predicates) crosslinguistically. Results show that the lengths of these word classes is significantly related to word order, that they can predict historical word order change in two different lineages, and that they account for more variance than family membership (descent) or areal patterns (geography) when examined using hierarchical linear regressions. Such findings suggest the possibility for a unified (“Min-Max”) theory of language evolution that can explain both observed surface structures in languages of the world as well as how such patterns have developed diachronically.
2 Complexity, processing, and word order
Linguistic complexity can take many forms, with many considerations regarding its measurement (see Jaeger & Tily 2011). As noted by Joseph and Newmayer (2012: 341):
“how linguistic complexity is conceived of, and how it might be identified and measured, has itself never been agreed upon by linguists. It cannot be neatly separated from theoretical considerations of where language structure ends and language processing (or use) begins, or where variation (whether social or stylistic) is located.”
Some current efforts to measure linguistic complexity focus on quantifying the overall complexity of a language (Bentz et al., 2016, 2017; Baerman et al., 2015, 2017), and such measures may be under considerable debate (see Bentz et al., 2023). Additional research suggests that word complexity can be measured by word length (see Lewis & Frank, 2016; and see Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022 for further discussion in relation to frequency in corpora). Words which are formed from longer sequences of phonemes are more difficult to process, and this is also related to their conceptual complexity. The fact that word length is related to a word’s complexity is important for our understanding of word order, i.e. how closely is the ordering of constituents related to how we process words/structures of varying lengths?
Research into how speakers process words has identified several constraints and tendencies that have various implications for how languages develop. Hawkins (2014) summarizes this literature in arguing for a comprehensive theory (the Performance-Grammar Correspondence Hypothesis; PGCH) based firmly in the idea of efficiency as an explanation for a large number of observed patterns and historical changes cross-linguistically. [^6] Levshina (2023) argues for another explanation termed “communicative” efficiency, which identifies “accessibility” as a factor in language use. This seems to overlap to some degree with the effects identified by “surprisal”. As Levshina’s observations do not make specific predictions about word order crosslinguistically, I revisit them in the discussion. His essential claim is that speakers minimize complexity (in forms, distance) in order to speed up processing, a claim that aligns with observations about constituent “weight” (Wasow, 2002; Zipf, 1949) and “dependency length minimization” (Futrell et al., 2015). However, this claim contrasts with findings from Surprisal Theory (Hale 2001; Levy 2008) that speakers maximize complexity, i.e. that more complex forms are preferred by speakers since they support information content/retrieval and reduce entropy in utterances (see Rezaii et al. 2023).
These two opposing perspectives on processing (the minimization vs maximization of complexity) make two different predictions for word order crosslinguistically. The PGCH claim (and Wasow’s) is that “heavier” or more complex constituents are typically “put off” until later in the sentence, while less complex constituents are put first. This predicts that a language in which nouns in general are easier to process than verbs should display SV order, while one in which verbs are easier to process should display VS order. Yet evidence from languages like Japanese, which places heavier phrases closer to the beginning of the sentence (Hawkins, 1994; Yamashita & Chang, 2001; Yamashita, 2002; and see discussion in Wasow, 2022) supports the opposing claim by Surprisal Theory that words with higher information content (or which improve predictability of following words) are more likely to occur first. Here the prediction is that a language with heavier nouns should display SV order, and one with heavier verbs should display VS order.
Such predictions, if true, should be observable in crosslinguistic corpora. An important note here is that while the processing literature is largely grounded in a functional/usage-based perspective, it does not necessarily require a usage-based stance. That is, cognitive processes can be argued to apply at various levels of processing, including above the “innate” bedrock of current Minimalist theory (see Rastelli 2025). If there is a universal processing bias in the brain related to the complexity of structures or words in language, as has been claimed, such a bias would apply regardless of whether grammatical rules are innate or learned, and would impact or drive how such rules are realized or change over time.
2.1 Main hypotheses
Using word length as a measure of word complexity, the current paper compares average lengths of particular classes of words (nouns/arguments and verbs/predicates) with word order in individual languages. The two processing claims in the literature give rise to the following two competing hypotheses regarding this relationship:
H1:
SV languages have shorter “S” than “V”, while VS languages have shorter “V” than “S”.
The average length of nouns is shorter in languages classed as Subject-initial, while the average length of verbs is shorter in languages classed as Verb-initial.
(minimize complexity)
H2:
SV languages have longer “S” than “V”, while VS languages have longer “V” than “S”.
The average length of nouns is longer in languages classed as Subject-initial, while the average length of verbs is longer in languages classed as Verb-initial.
(maximize complexity)
2.2 Investigating word order and length
Using the 1,597 languages in the tagged PBC, we can identify basic word order via the N1 ratio. [^7] As an example, English [ISO 639-3: eng] has an N1 ratio of 4.51 and is classed as SV, whereas Irish [ISO 639-3: gle] has an N1 ratio of 1.24 and is classed as VS. Details on the N1 ratio and its validity are given above and at the linked repository (Fn 4). We can also identify the average length of nouns/arguments and verbs/predicates in each language’s corpus and see how well this correlates with the N1 ratio.
2.2.1 Word order
Of the languages in the tagged PBC, 976 are also found in typological databases with categorizations for word order. For the remainder, the work described at the linked repository has shown that the N1 ratio is a valid corpus-based proxy for word order. Accordingly I used the N1 ratio to impute the (SV, VS, free) category of the remaining 621 languages in the tagged PBC not categorized by WALS, AUTOTYP or Grambank. [^8] I trained a Gaussian Naive Bayes classifier on the known data to predict the category (SV, VS, free) of languages for which word order status was unknown, based on the N1 ratio as derived from the PBC. Code to reproduce these results can be found at the linked repository (Fn 4). This resulted in 1,251 languages classified as “SV”, 295 languages classified as “VS”, and 51 languages classified as “free”.
2.2.2 Nouns and verbs
To measure word length, I counted the number of characters in a word tagged as Noun/Argument or Verb/Predicate and computed an average length for each class on a per-language basis. [^9] For example, English [ISO 639-3: eng] arguments have an average length of 5.9 characters, while English predicates have an average length of 5.75. Irish [ISO 639-3: gle] arguments have an average length of 6.31 characters, while Irish predicates have an average length of 6.29 characters. In the tagged PBC corpus the average number of (unique) arguments per language is 732, while the average number of predicates is 495. The mean length of arguments is 6.76 characters, while the mean length of predicates is 6.90 characters. Conducting a Student’s T-Test showed that this is statistically significant (p = 0.006), indicating that crosslinguistically, arguments are shorter than predicates. All statistics showed a bimodal (normal) distribution.
2.2.3 Relationship between word order and noun/verb lengths
To see whether word order is related to the length of nouns and verbs I conducted a repeated-measures ANOVA with the average lengths of nouns and verbs as the repeated measures and the word order (SV, VS, free) as the between subject factors. Analysis of the results of word order and its relation to length of the different argument and predicate categories is quite striking. First, both SV and VS languages have a highly significant effect of word length, [^10] Additional statistical comparisons and plots are available at the linked repository (Fn 4). whereby arguments are, overall, shorter than predicates crosslinguistically (Figure 1; p < 0.001). [^11] It is also worth noting here that in this corpus words are overall shorter in SV languages than VS languages, which in turn are shorter than those in ‘free’ languages. This displays a general observed trend that languages with free word order tend to have more morphology to identify argument roles and other grammatical relations (Koplenig et al. 2017; Yadav et al. 2020).
Figure 1: Normalized lens of Args, Preds, by word order
Figure 2: Frequency-weighted lens of Ns, Vs, by word order
If we consider only nouns and only verbs, the findings are similar when we normalize the dataset by reducing the vocabulary to unique items. However, if we allow for frequency effects (i.e. we do not reduce the vocabulary to unique items) there is a clear correlation with word order: nouns are longer than verbs in SV languages, but shorter than verbs in VS languages (Figure 2; p < 0.001). This is a major finding that we will return to in the discussion. Setting it aside for now, we turn to two different tests of this measure relating to diachrony: predicting word order change, and the potential interaction with family membership (descent) and area (geography).
2.3 Testing the predictive validity of word length
Given the observation that relative lengths of nouns and verbs differ in corpora for languages with different word orders, we can ask how predictive this observation is. That is, can we predict the word order of a language based on the relative lengths of its nouns and verbs? If we are strictly interested in identifying basic word order from corpora, it is clear that the N1 ratio already allows us to do so. However, we are also interested in understanding how noun and verb lengths may be related to word order. Being able to predict word order from noun and verb lengths in a corpus for both a modern language and its ancestral language would indicate that such properties are related to how word order evolves across time.
To investigate this question, I sourced additional corpora of tagged data from two pairs of historical and modern languages in which the modern language changed its word order: 1) Ancient Hebrew (VS) and Modern Hebrew (SV) and 2) Classical Arabic (VS) and Egyptian Arabic (SV). All languages belong to the Semitic branch of the Afroasiatic language family, a necessary constraint given the limited amount of data available for most pre-modern languages. In the case of Hebrew, the UDT has POS-tagged sentences for both Ancient [^12] https://universaldependencies.org/treebanks/hbo_ptnk and Modern [^13] https://universaldependencies.org/treebanks/he_htb varieties. In the case of Classical Arabic I used the Quranic Arabic Corpus [^14] Version 0.4: https://corpus.quran.com and for Egyptian Arabic I used the BOLT Egyptian Arabic Treebank (Maamouri et al. 2018).151515(Discussion Forum) https://catalog.ldc.upenn.edu/LDC2018T23 If the lengths of nouns and length of verbs in these corpora allows us to correctly classify Ancient Hebrew and Classical Arabic as VS and the modern varieties as SV, this would suggest that these are salient properties affecting how word order has developed diachronically. Importantly, given the textual differences between the tagged PBC (New Testament texts) and the test corpora (Biblical, Quranic, modern, and informal texts), such a finding would be significant.
In order to conduct this test, the data for Ancient Hebrew (ISO 639-3: hbo), Modern Hebrew (ISO 639-3: heb), Classical Arabic (cla) [^16] Classical Arabic does not have an assigned ISO 639-3 code or Glottocode. Accordingly, I use the abbreviation cla here, as there is no current language assigned to this code in the ISO 639-3 tables. and Egyptian Arabic (ISO 639-3: arz) were processed and the respective word classes were counted and measured as per the method outlined for the tagged PBC (each corpus > 400 sentences, see Table 1). The average length of nouns and verbs (including frequency effects) were then used as continuous features for a Gaussian Naive Bayes (GNB) classifier.
Language ISO 639-3 Data Source Sentences Unique nouns Unique verbs
Ancient Hebrew hbo UDT v2.14 409 734 782
Modern Hebrew heb UDT v2.14 882 2068 1297
Classical Arabic cla Quran 6,236 8,243 7,966
Egyptian Arabic arz BOLT 31,688 16,698 10,581
Table 1: Corpora used for testing historical word order change
Three additional features were also used to expand the feature dimensions. The first is a categorical feature: whether a language’s nouns are longer than its verbs. The second is a continuous feature: the difference in length between nouns and verbs. This second measure is related to minimization of complexity, based on the idea that a larger difference between lengths of word classes may effect the likelihood of that particular class being placed first. The third measure is the ratio of noun and verb lengths, and is related to maximization of complexity (via “entropy”), whereby a lower ratio indicates that the particular word classes are less likely to be easily differentiable (and therefore predictable) based on their length.
Using these five features, a GNB classifier was trained on the original combined word order dataset containing hand-coded classification of 974 languages (Modern Hebrew and Modern Arabic were removed), and then used to predict word order for our pairs of languages: Ancient Hebrew and Modern Hebrew on the one hand, Classical Arabic and Egyptian Arabic on the other. Ancient Hebrew was classified as VS while Modern Hebrew was classified as SV by the trained classifier. Classical Arabic was classified as VS while Egyptian Arabic was classified as SV. This indicates that the relative lengths of nouns and verbs, when frequency is taken into account, can predict word order from corpora, with various implications for our understanding of language processing and language change. This is a striking finding, and I am unaware of any other research that has been able to show predictive validity of a model related to word order change. However, it is possible that this result is conditioned to some degree by the fact that these are Semitic languages spoken in the same geographical area, so we now turn to an investigation of whether descent or location provides a better explanation for the correlation of noun and verb lengths with word order.
2.4 Testing variance predicted by word length
Dunn et al. (2011) showed that descent from a common ancestor was a better predictor of many grammatical patterns than other explanations. Although Jäger & Wahle (2021) counter this finding with additional data, we can also assess this claim using the current dataset to see if descent is more explanatory of word order for a given language than the lengths of nouns and verbs. Additionally, in more recent work, Naranjo & Becker (2022) show that areal affects are as important as family membership in predicting linguistic phenomena, finding that the results in Dunn et al can be just as easily predicted by language area. This is also a claim that we can test with our dataset. Accordingly, this investigation can proceed with several (motivated) simplifying observations.
First, for each language we can take the highest level node of family membership that is known based on current scholarship (via Glottolog). [^17] Glottolog (Hammarström et al. 2024; https://www.glottolog.org) contains expert annotations for language family affiliations for over 7,000 languages, along with additional aggregated data from a variety of sources. This is because all modern languages are the same age, and while their pathways/histories may differ, this does not impact whether they can be considered part of the same phylogenetic tree. That is, if descent is more explanatory than word class length, we would expect languages belonging to the same tree to have a higher likelihood of sharing the same word order than not. Second, we can use the macroareas identified by Glottolog as well as the actual GPS coordinates of languages to control for areal affects. Third, we can run various analyses with differing sets of data in order to see how this affects the statistical robustness of the findings. It may be, for example, that using a smaller sample gives a different result than a larger or more diverse one. Here, the more languages we have for a language family or geographical location, the higher the likelihood that any relationship identified is not due to chance. Finally, we can remove languages with “free” word order from consideration, as this leaves us with a binary distinction in word order (SV, VS) that opens up more possibilities for statistical tools.
In particular, a binary dependent variable for word order means we can use hierarchical linear regression, which allows us to observe the relative impact of particular features that are theoretically relevant to the question at hand by seeing how they influence the performance of successive models. In the case of word order, we can train a regression model with a feature known to be relevant to word order (the N1 ratio), and then subsequent models that incorporate additional features of interest (“language area/location”, “family membership”, “noun/verb lengths”). We can then compare the models to see whether the additional features increase the variance explained in the data via the F-value: an increase in F-value (change) and the significance of this change indicates more/additional variance explained by the particular feature added in a model.
Accordingly, for all languages in the tagged PBC I retrieved the family membership and geographical information from Glottolog, filling in missing data via Ethnologue and other sources. Then I filtered the data on four conditions: a) languages found in Dunn et al that are also present in the tagged PBC (106 languages), b) languages found in the tagged PBC that are in the four families investigated by Dunn et al (539 languages), c) languages in families with more than 75 members in the tagged PBC (962 languages), and d) languages in families with 2 or more members in the tagged PBC (1,486 languages). For each of these conditions, I ran a hierarchical linear regression, adding features of interest progressively to see how they impacted the fit of the successive models and whether particular features account for additional variance.
The base model (Model 1) uses the N1 ratio as the primary feature for classification. This is because we know from comparison with expert determinations of word order (see footnote 5 above) that the N1 ratio significantly correlates with the classification of languages as SV/VS. Due to the strong relationship observed between language area and typological properties of language (see Naranjo & Becker 2022), Model 2 then includes GPS coordinates and macroarea identifiers for each language as features. Model 3 additionally incorporates descent (language family membership) as a feature (as per Dunn et al. 2011), and Model 4 includes the average lengths of nouns and verbs (as per the relationship identified in this paper). If descent is more explanatory than the lengths of nouns/verbs, this should be clear from the differences between the respective models for each condition. Results are displayed in Table 2.
a) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 106 45.85 7.8e-10
2 (Lg area) 106 20.40 2.4e-12 8.58 3.9e-05**
3 (Family) 106 16.70 5.8e-12 1.48 0.227
4 (N/V len) 106 13.40 4.9e-12 3.27 0.042*
b) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 539 230.41 1.4e-43
2 (Lg area) 539 100.00 1.7e-63 39.86 2.9e-23**
3 (Family) 539 80.71 5.3e-63 2.46 0.118
4 (N/V len) 539 64.75 3.5e-67 14.57 6.9e-07**
c) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 962 426.74 9.9e-79
2 (Lg area) 962 220.49 4.2e-134 105.35 6.1e-59**
3 (Family) 962 176.22 6.6e-133 0.04 0.839
4 (N/V len) 962 130.93 8.3e-135 9.69 6.7e-05**
d) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 1486 586.46 1.8e-109
2 (Lg area) 1486 193.61 1.3e-133 45.19 6.0e-28**
3 (Family) 1486 155.10 1.2e-132 1.03 0.309
4 (N/V len) 1486 115.43 3.5e-135 11.02 1.7e-05**
Table 2: Hierarchical Linear Regressions under 4 conditions
From these analyses we can see that, as expected, the N1 ratio significantly predicts word order under all conditions. Additionally, language area is a significant predictor of word order that adds to the strength of the classification. Language family (descent) does not add additional predictive power to our model under any conditions (over language area), whereas the length of nouns and verbs does. This analysis shows that the length of nouns and verbs in a language accounts for more variance and at the same time is more significant than that language’s family membership in determining its word order classification, under all test conditions. When we remove language area from the linear regression this is still the case, [^18] See additional statistical results at the linked repository (Fn 4). indicating that while language area is a stronger predictor than family membership, lengths of nouns and verbs is yet a stronger one.
This largely aligns with Jäger and Wahl’s (2021) finding regarding the (lower) correlation between descent and linguistic structures in comparison to other explanations, with the caveat that Jäger and Wahl included language isolates, which is likely to greatly diminish the affect of descent. If a large number of languages in a sample are not related to others, clearly descent can have only a minimal impact on word order. By including only language families with two or more members (which excludes 111 languages in the taggedPBC), I show that descent does not explain much variance above and beyond language location/area, in contrast with noun and verb lengths. This also suggests that language family/descent is highly correlated with linguistic area, which conforms with expectations related to human history and migration. Additionally, the amount of variance explained seems to depend on the number of languages (and language families) in the sample, highlighting the need for more data for languages from under-represented language families.
3 Discussion
That arguments overall are shorter in languages of the tagged PBC, across all word orders, suggests that there are certain (non-Noun) arguments affecting our observations. Particularly, the effect of PRON seems quite likely, whereby a tendency for pronouns to be shorter crosslinguistically, along with a preference for pronouns as arguments, may impact the observation that arguments are shorter overall than predicates. This pattern aligns with the processing claim that ‘heavier’ words (or word classes in this case) occur later in a sentence, partly validating Hypothesis 1. The counter-argument, also from the data, is that nouns are longer than verbs in SV languages and shorter than verbs in VS languages when frequency is considered, validating Hypothesis 2. This suggests that speakers prefer to place heavier (more “contentful”, or perhaps “novel”) constituents before lighter ones crosslinguistically in many cases.
These seemingly antithetical results partially validate both the “weight”-based theory and the “surprisal” theory, highlighting the interaction of two pressures: one whereby speakers prioritize lighter constituents (possibly for the purposes of production/processing efficiency), and one where they prioritize heavier constituents (possibly for informational purposes). The two perspectives may in fact be looking at the same data from different angles. On the one hand, grouping all arguments into a single class highlights the overall impact of shorter strings, since the most frequently used arguments (pronouns, demonstratives) are typically shorter crosslinguistically. On the other, more “content-ful” word types that aid planning and inference or predictability (nouns, verbs) will typically be longer crosslinguistically.
These two pressures can be (and, in fact, have been) considered as two different kinds of “efficiency”: production efficiency vs communicative efficiency. Under this view, production efficiency is related to length of words (as per Wasow 2002, 2022; Hawkins 2014) - speakers use shorter words whenever they can to efficiently produce utterances (see also Mahowald et al. 2013). The counter-pressure, communicative efficiency, is related to making a speaker’s utterances more readily accessible/interpretable by recipients (as per Levshina & Moran 2021; Levshina 2023) - speakers prioritize words that contain more meaningful, informative content in order to aid communication. Once context has been established, these longer words can be replaced by more efficient (shorter) terms such as demonstratives or pronouns. [^19] For example, there is a common crosslinguistic pattern of using a noun first to establish reference, then switching to pronouns, i.e. "that tree, it is big, it will fall…" See also discussion of the (lack of) encoding of subjects crosslinguistically in Berdicevskis et al. (2020).
I would argue, however, that conceptualizing “efficiency” in these two distinct ways, as currently done in the literature, runs the risk of confusing the two. It is in fact not efficient (from a processing perspective) to maximize information, as this also tends to maximize complexity. Instead, I propose a different framing of these principles.
3.1 A “Min-Max” theory of processing
The current proposal re-formulates or refines Hawkins’ PGCH to incorporate (increased) complexity as a motivation for the order of constituents, bearing some resemblance to proposals in Levshina (2023). I frame this as (to borrow a term from computer gaming) the “Min-Max” theory of processing (schematized in Figure 3). The idea here is that speakers seek to minimize their processing effort while maximizing the information in their utterances. These are competing/parallel pressures that speakers are constantly balancing in production (or “performance”) in order to choose the best strategy for their immediate goals. As speakers produce variants along a cline of grammaticality, these variants compete on the basis of multiple factors, including how well such variants support Min-Max-ing in relation to other variants, leading to language change.
Figure 3: Min-Max schema
These two competing pressures offer a more complete (and parsimonious) explanation for word order change (and possibly other kinds of linguistic change) than existing theories of language. In one possible scenario the frequent use of short pronominal forms encourages drift toward an argument-initial order, and nouns subsequently lengthen as they provide more information to speakers and move toward an initial position. This, combined with the preponderance of shorter pronominal forms crosslinguistically, may provide an explanation for why the majority of languages reported in the world are SV. Conversely, if a language has or develops longer verbs, the preference for these heavier constituents to occur first could put pressure on the language to maintain or develop predicate-initial order, which would explain why some VS languages are quite stable. In each case the direction of development would depend on a number of factors that have been implicated in discussions of language change and grammaticalization (see Heine & Kuteva 2002; Matras 2011; Narrog & van der Auwera 2011). However, the effect of word length is clearly a major part of the explanation for why we can observe (and predict) such a difference in word order between Ancient and Modern Hebrew as well as Classical and Egyptian Arabic via their respective corpora.
The existence of multiple competing pressures affecting the development of language, as described by this Min-Max theory, is supported by work in clinical psychology (e.g. Rezaii et al. 2023) and language modeling (e.g. Hahn & Xu 2022), as well as findings from information theory (Zaslavsky 2020; Tucker et al. 2025). In a study of patients with nonfluent aphasia, Rezaii et al. (2023) showed that patients with this condition compensate for their deficit in producing complex grammatical structures by increasing the information content of their utterances, dropping high-frequency (function) words in favor of low-frequency (content) words. By modeling variations in grammatical structure based on corpora from 80 languages, Hahn & Xu (2022) found that the resulting models were less optimized for dependency length and information locality than the original structures of each language, indicating that languages evolve to efficiently balance the two considerations. [^20] Notably, Hahn & Xu propose a theory that the interaction between Dependency Length (DL) and Information Locality (IL) accounts for word order variation and change, but neither of the two concepts, as expressed, make claims about directionality/order, only about the relationships between (internal) constituents. In contrast, the Min-Max theory makes the claim that speakers balance information content (as reflected by longer nouns in SV languages and longer verbs in VS languages) with processing/production concerns, and that this interaction is what drives the evolution of word order. As nouns lengthen or verbs shorten in a language, this will affect their placement in a stream, which then impacts the overall order observed in surface structures. That is, the Min-Max theory makes a universal claim regarding language-external pressures, while DL and IL are language-internal concepts. These and other studies indicate that language users are sensitive to efficiency as it pertains both to processing and information content.
The sensitivity of language users to processing efficiency and information content has also been a major topic of investigation in information theory, which is concerned in part with the encoding and decoding of information streams. One information-theoretic proposal is that semantic systems evolve following an “Information Bottleneck” principle (IB; Tishby et al. 1999), whereby “optimal representations satisfy a tradeoff between compressing X, i.e. minimizing the representational complexity, and maximizing the relevant information about Y” (Zaslavsky 2020: 14). Work by Zaslavsky shows that IB can account for the development of color naming systems in language among other kinds of sytems, and additional research shows the utility of IB in machine learning and distributional models of semantics (Slonim & Tishby 2000).
The implications of this principle have most recently been examined in agent-based emergent learning simulations with the addition of a “utility” principle (Tucker et al. 2025) – the interaction between “utility”, “informativeness”, and “complexity” allows more human-like communication to emerge among agent models in their study. Such a finding is very much supported by the results reported here, though I would argue that the Min-Max theory offers a more parsimonious account of human behavior as related to language specifically. While the current corpus evidence applies at the word level in relation to basic word order, it seems likely that the minimization of processing difficulty would interact with the maximization of information content at multiple (hierarchical) levels. Speakers min-max in order to select words in a stream (“informativeness” and “complexity” as per Tucker et al.), while simultaneously min-maxing to select relevant structures related to their communicative goals (Tucker et al.’s “utility”).
The Min-Max theory also bears some resemblance to Levshina’s (2023) proposals, described in terms of communicative efficiency and accessibility. However, such indicators cannot fully explain certain patterns in grammar such as “harmonic” word order. A recent proposed explanation for such patterns involves frequency of word classes, whereby Mansfield & Krapp (2025) suggest that “harmonic ordering is really a frequency effect: as replication converges on a phrase structure with a consistent linear order, the most frequent word class tends to be at one edge.” Although they did not test token length in their model, it is notable that in the current study frequency is a major factor in identifying the relationship between length and word order, suggesting that the most frequent word class in their model may also be the longest element of the domain they are testing, which would be driven by a maximization of information principle. The Min-Max theory also seems to provide an explanation for non-parallel (hierarchical) harmonic word order, which is not explained by their model.
The Min-Max theory may additionally provide an explanation for why word forms remain long (or lengthen), as described for particular languages (see Cristofaro 2021; Haspelmath 2021; Petré 2017). Here, again, the maximization of information principle would support the development or retention of particular (long) word forms, as speakers balance this need against the minimization of processing difficulty. As noted above, this proposal is supported by evidence from color-naming systems, which show patterns in keeping with the Information Bottleneck principle. The Min-Max theory would suggest that this principle does not only hold for “forms” (of, i.e. color words), but also governs “behavior”: speakers behave in ways that are consistent with maximizing information in streams (via word order), while minimizing processing difficulty.
As noted in section 2, such a theory can be conceived of as a universal (underlying) pressure that applies generally to how language is realized. From an innate perspective, speaker choices impacted by such pressures would lead to a re-analysis or re-formation of rules that guide surface realizations. From a usage-based perspective such pressures would lead to the development of new surface variants. These are not necessarily opposing views (see Rastelli 2025), and the outcomes are quite similarly dependent on language being viewed as a complex adaptive system, whereby diachronic changes may have a combination of both “result-oriented” and “source-oriented” explanations (see Cristofaro 2019; Schmidtke-Bode & Grossman 2019).
Another observation from the results reported here is that information content seems to trump length. We find that long nouns are used with more frequency in SV languages, while long verbs are used with more frequency in VS languages. This indicates that (maximizing) information content is more important to speakers than (minimizing) processing complexity. This could explain the “agent/subject-first preference” in experiments (Sauppe et al. 2023; Culbertson et al. 2012), as the “information maximization” pressure wins out in cases where this most efficiently meets communicative goals.
Further, the data shows that all languages have a relatively high proportion of sentences in which the argument occurs before the predicate. In VS languages, the proportion of argument-initial to predicate-initial sentences is 1/2 or more (min: 0.56 n/v, 0.65 a/p), highlighting that even VS languages still place the noun/argument first in many constructions. Similarly, while all SV languages have a high proportion of N1 sentences (max 2.43 n/v, 7.17 a/p), the number of V1 constructions is non-zero, supporting a gradient approach to word order (Levshina et al. 2023).
The fact that the length of particular word classes can predict the basic word order of a language is a novel finding, leading to some general predictions: if, on average, nouns are longer than verbs in a language, it is highly likely to have SV word order. If, on the other hand, verbs are longer, it is highly likely to have VS word order. The effect of frequency cannot be ignored either, highlighting the importance of conducting large-scale corpus based studies to investigate this question. Since languages are used by speakers and such speakers use certain words more than others, we can only observe this frequency effect (or “weight”) of word class length via corpora.
The idea that speakers balance multiple considerations in production is not new (Pike 1967; Jackendoff & Pinker 2005; Seržant & Moroz 2022; Hörberg & Sjons 2023) but what is novel is the idea that all humans experience a specific universal pressure to balance production/processing efficiency with information content. This pressure is observable crosslinguistically, and would apply at multiple processing levels, regardless of whether a speaker is accessing a universal grammar (“innate”) or inferring structure on the basis of the input they receive (“usage”). Following on from Zipf’s and subsequent observations, there is no doubt that the length of strings is an important factor in processing streams, but clearly not all strings are created or treated equally by speakers. This study shows that nouns are treated differently (or at least have a different distributional relationship) than other kinds of arguments, which leads to different realizations of word order crosslinguistically. The pressures identified here have wide-ranging implications and applications for any conceptualization of how language operates, whether in terms of understanding the intricacies of individual linguistic scenarios, designing intelligent systems that use language, or developing clinical interventions to assist speakers in various ways.
3.2 Limitations
One limitation in this study is the use of automated processes to identify part of speech tags in the data. Ideally, all the data should be hand-annotated for parts of speech, which would allow for better comparison of the different word categories across languages. Given the time consuming nature of this task and the need for expert annotators, this is not yet possible, but ongoing research could address this by sourcing experts to annotate the data. Although I have shown (in Ring 2025 and via the linked repository) that the translation/alignment method does correspond well to hand-tagged data, it may be that this could be improved by tagging a small amount of data for each language in the PBC in order to train POS taggers that perform better than the current translation method.
Another limitation is that the dataset used, while exceedingly large in comparison to previous efforts, still only represents about a quarter of the world’s languages. Since the majority of the world’s languages are not represented here, it is not clear that these findings can be fully generalized to all languages of the world, though I have shown in section 2.4 that much can be learned from sampling. Additionally, the contrast between intransitive SV/VS/free word order is somewhat of an oversimplification of word order patterns. Given the current obstacles to annotation, it is a necessary simplification for this study, however future work could expand to a more diverse set of word/constituent orders and potential relationships with word length and/or complexity. Given the nature of the dataset used, it is also possible to pursue a gradient approach to word order, as advocated by Levshina et al. (2023). The issue of word complexity also could use further expansion to cover a broader range of possibilities than simply the length of a word based on a set of (roman) characters.
A final note is that our conclusions are only as good as the data that we have available to us. The tagged PBC is a significant step in the right direction, toward a representative sample of large tagged corpora for languages of the world. Systematic investigation of this dataset can give us insight into language typology (the current state of language), as well as pressures that impact this realization in diachrony, both in terms of how languages have developed (their history) and how they might develop (their future). On the basis of this investigation I have proposed a parsimonious “Min-Max” theory of processing and information structure that underlies language evolution, whereby speakers minimize processing effort while maximizing information gain. This proposal is in keeping with at least one existing information-theoretic proposal (IB) as well as theories of processing (PGCH; surprisal) and can explain the correlation between the length of nouns/verbs and basic word order, given that these features allow us to predict the word order of a language. It is as yet unclear whether such a principle would apply at multiple hierarchical levels within language, which is an area that deserves further careful investigation with datasets such as the tagged PBC. For now, I have demonstrated that noun/verb lengths are more explanatory than descent or geography and that for languages with corpus data, word length predicts word order.