### v2

# Word length predicts word order: “Min-max”-ing drives language evolution

Abstract
A fundamental concern in linguistics has been to understand how languages change, such as in relation to word order. Since the order of words in a sentence (i.e. the relative placement of Subject, Object, and Verb) is readily identifiable in most languages, this has been a productive field of study for decades (see Greenberg 1963; Dryer 2007; Hawkins 2014). However, a language’s word order can change over time, with competing explanations for such changes (Carnie & Guilfoyle 2000; Crisma & Longobardi 2009; Martins & Cardoso 2018; Dunn et al. 2011; Jäger & Wahle 2021). This paper proposes a general universal explanation for word order change based on a theory of communicative interaction (the Min-Max theory of language behavior) in which agents seek to minimize effort while maximizing information. Such an account unifies opposing findings from language processing (Piantadosi et al. 2011; Wasow 2022; Levy 2008) that make different predictions about how word order should be realized crosslinguistically. The marriage of both “efficiency” and “surprisal” approaches under the Min-Max theory is justified with evidence from a massive dataset of 1,942 language corpora tagged for parts of speech (Ring 2025), in which average lengths of particular word classes correlates with word order, allowing for prediction of basic word order from diverse corpora. The general universal pressure of word class length in corpora is shown to give a stronger explanation for word order realization than either genealogical or areal factors, highlighting the importance of language corpora for investigating such questions.
1 Introduction
There are two major approaches to studying word order. The first (typological) approach is to observe word orders in natural languages and identify patterns whereby the orders of certain constituents are correlated with other grammatical patterns. This approach has developed multiple crosslinguistic databases with expert determinations of word order (WALS, Dryer & Haspelmath 2013; AUTOTYP, Bickel et al. 2023; Grambank, Skirgård et al. 2023) and has identified statistical universals in which the orders of certain elements increase the likelihood for other patterns to be present in a language (Greenberg 1963; Dryer 2007, 1992, 2011; Verkerk et al. 2025). The second (processing) approach focuses on experimental evidence for how constituents in particular languages are processed in the brain. This approach has identified potential universal tendencies or preferences in processing, notably a preference for “agents first” (Culbertson et al. 2012; Futrell et al. 2015a; Sauppe et al. 2023), “efficiency” (Zipf, 1949; Wasow, 2002; Futrell et al., 2015b), and “surprisal” (Hale 2001; Levy 2008).
While the typological observations do not make predictions about a language’s basic word order (the universals take basic word orders as a starting point and make further predictions, see Verkerk et al. 2025), the processing research does imply some particular claims and predictions about how words should be ordered in languages of the world. For example, the “agents first” (or “Subject first”) bias found in experiments suggests that there is a universal psychological pressure for speakers of all languages to produce agent arguments (“do-ers” of actions, typically represented by animate nouns and pronouns) before other kinds of constituents in sentences. This aligns with observations of the dominant crosslinguistic word order being Verb-final or Subject-initial (Schouwstra & de Swart 2014; Schouwstra et al. 2022).
At the same time, findings from language processing studies suggest that other pressures mediate the realization of word order patterns, namely in relation to “efficiency” and “surprisal”. The efficiency literature largely follows from Zipf’s (1932; 1949) observations that a) languages show a logarithmic distribution whereby the most frequently used words occur much more often than the next most frequent, and that b) the most frequent words in a language are also the shortest. The observation that shorter (or less complex) elements are more likely to occur initially in a sentence, while longer (or more complex) elements occur later, can be traced to Behaghel (1909), and the general idea is that this supports efficient communication. In contrast to this view, the surprisal literature focuses on how salient or predictable items are, suggesting that longer (or more complex) words are more likely to occur first in a sentence since they improve predictability of following constituents (Hale 2001; Levy 2008; Sauppe et al. 2021). This is supported by clinical evidence (Rezaii et al. 2023) and information theoretic studies (Zaslavsky 2020; Tucker et al. 2025).
Such studies have implications for particular languages regarding the ordering of constituents. For example, if a language has longer verbs than nouns, we would expect the language to be Verb-final based on the efficiency literature (longer or more complex words last), whereas the surprisal literature would predict it to be Verb-initial (longer or more complex words first). These competing predictions represent hypotheses that have yet to be tested to any great degree (though some attempts have been made, at least in relation to dependency structures, see Jing et al. 2021; Hahn & Xu 2022).
The lack of testing highlights the complexity of investigating pressures that influence the evolution and realization of language structures, as well as the dearth of datasets and databases that could address these questions. [^1] Datasets that have been used by previous research include the Universal Dependencies Treebanks (Zeman et al. 2024) and ODIN (Xia et al. 2016) projects. The UDT contains 160+ hand-annotated languages available for crosslinguistic investigation, but these are not parallel texts and the number of sentences in each language is variable, with some having fewer than 200 sentences in the v2.14 release. The ODIN project (and its re-packaged form via LECS: https://huggingface.co/datasets/lecslab/glosslm-corpus) has interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences. Additional cross-linguistic investigations using tagged corpora include Hahn et al. (2020: 51 languages), Jing et al. (2021: 71 languages), and Hahn et al. (2022: 80 languages). For each of these studies the majority of languages are from the Indo-European language family. Besides limiting generalizability of findings, such data risks bias toward a particular group of languages (Blasi et al. 2022), in much the same way that other scientific findings have historically been biased toward WEIRD demographics (Henrich et al. 2010). In particular, questions regarding the length of constituents in relation to their order in sentences requires a large number of sentences tagged for parts of speech (at minimum). Given that languages develop over time via cultural evolution, it is also difficult to separate general cognitive pressures from historical (genealogical) or areal factors. This means that inferring crosslinguistic or universal principles necessitates both a large number of languages as well as sufficient diversity of languages to establish statistical validity.
Additionally, the fact that languages develop over time through communicative interaction requires that any explanation for universal principles be related to such interactions. That is, in an evolutionary linguistic paradigm it is individuals (speakers) in a population (language) that drive variation and select particular linguistic variants for propagation. The selection process itself is likely motivated by various factors, two of which are combined in the Min-Max theory of language behavior. The Min-Max theory proposes that in production speakers select variants based on two competing pressures: the pressure to minimize effort and the pressure to maximize information (Figure 1). This language behavior is considered to be core to what drives selection and propagation of linguistic constructions, and thus change over time (in, i.e. word order).
Figure 1: Min-Max schema
The current paper takes advantage of a massive tagged parallel corpus to show how these pressures are realized in actual language data in relation to word order. While the dataset is still under development, the baseline data is tagged for parts of speech, which allows us to derive a proxy measure for word order from individual language corpora, as well as to measure other linguistic properties such as the average length of various word types. Results show that basic word order in intransitive clauses is highly correlated with the length of particular word classes, allowing for the prediction of a language’s word order based on corpora not in the dataset and explaining variance better than both genealogy and areal factors.
In the remainder of this paper I first highlight the properties of the dataset (§2). I then discuss three studies that investigate word length in relation to word order (§2.1), the predictive validity of word lengths for word order (§2.2), and the variance predicted by word length in regression models (§2.3). I conclude with a discussion of the implications of these studies in relation to language evolution (§3), presenting the Min-Max theory of language behavior (§3.1) as well as some limitations and future directions regarding the current effort (§3.2).
2 Methodology and hypotheses
The taggedPBC (Ring 2025) [^2] https://github.com/lingdoc/taggedPBC - see fn 5 for a link to a minimal OSF repository that allows for replication of the studies in this paper. contains pos-tagged parallel text data in 1,942 languages (roughly 25% of the world’s languages), with over 1,800 parallel verses for more than 1,850 of those languages. The dataset represents 155 language families and 78 isolates, dwarfing previously available resources, and the tagging methodology (via crosslingual tag transfer) has been shown to correspond well with existing automated taggers for high-resource languages (as found in the SpaCy and TranKit NLP libraries). [^3] https://spacy.io ; https://nlp.uoregon.edu/trankit Additionally, a novel measure extracted from individual corpora in this dataset (the “N1 ratio”) is significantly correlated (p < 0.001) with expert determinations of a language’s basic word order for intransitive clauses (“SV”, “VS”, “free”; Figure 2) as identified in the typological databases WALS (Dryer & Haspelmath 2013), Grambank (Skirgård et al. 2023), and Autotyp (Bickel et al. 2023). [^4] The procedure for determing the N1 ratio is essentially to count the number of sentences in a corpus that have both arguments (nouns, pronouns, etc) and predicates (verbs, auxiliaries), and then to derive a ratio from the number of argument-first sentences divided by the number of predicate-first sentences in that set. As an example, English [ISO 639-3: eng] has an N1 ratio of 4.51, whereas Irish [ISO 639-3: gle] has an N1 ratio of 1.24, based on their respective corpus in the taggedPBC. This ratio can then be related to other aspects of language, such as word order classifications. See Ring (2025) for further details. What this means is that using this dataset we can derive word order and word length information in a systematic manner, allowing for an investigation of word length in relation to word order.
Figure 2: The N1 ratio and word order in 3 typological databases (p < 0.001; Ring 2025)
The following studies focus on intransitive word order (i.e. the basic word order of intransitive clauses, in which a predicate requires a single “S” argument) in languages found in the taggedPBC. [^5] Minimal data and Python code to replicate the results in this paper can be found at https://osf.io/nsaq2 (3MB). More comprehensive data and code (7GB) can be found at the taggedPBC github repository (fn 2). The first study measures word classes, showing that average lengths of nouns and verbs in a language’s corpus are highly correlated with SV, VS, and free word order classifications when frequency is considered. The second study investigates whether this effect occurs outside of the taggedPBC, showing that information on average lengths of nouns and verbs computed from different corpora also allow for the classification of languages as SV or VS, differentiating between historical and modern languages for which word order changes are known. The third study examines whether the lengths of nouns and verbs in corpora provides a better indication of word order than known areal or genealogical signals, and shows that in heirarchical linear regressions, word length is indeed a better predictor of word order.
2.1 Study 1: Word length and word order
Study 1 follows previous studies in using word length as a measure of word complexity (Lewis & Frank 2016; Bentz et al. 2016, 2017; Baerman et al. 2015, 2017; Bentz et al. 2023; and see Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022 for further discussion in relation to frequency in corpora), comparing average lengths of particular classes of words (nouns/arguments and verbs/predicates) with word order in individual languages. The two processing claims in the literature (from “efficiency” and “surprisal” approaches) give rise to the following two competing hypotheses regarding this relationship:
H1:
SV languages have shorter “S” than “V”, while VS languages have shorter “V” than “S”.
The average length of nouns is shorter in languages classed as Subject-initial, while the average length of verbs is shorter in languages classed as Verb-initial.
(minimize complexity)
H2:
SV languages have longer “S” than “V”, while VS languages have longer “V” than “S”.
The average length of nouns is longer in languages classed as Subject-initial, while the average length of verbs is longer in languages classed as Verb-initial.
(maximize complexity)
For the 1,942 languages in the taggedPBC, we can identify basic word order via the N1 ratio. We can also identify the average length of nouns/arguments and verbs/predicates in each language’s corpus and see how well this correlates with the N1 ratio.
2.1.1 Word order
Of the languages in the taggedPBC, 1,152 are also found in typological databases with categorizations for word order. For the remainder, the N1 ratio was used to impute the (SV, VS, free) category of the remaining 790 languages in the taggedPBC not categorized by WALS, AUTOTYP or Grambank, and the final categorizations were stored for all 1,942 languages. [^6] I trained a Gaussian Naive Bayes classifier on the known data to predict the category (SV, VS, free) of languages for which word order status was unknown, based on the N1 ratio as derived from the taggedPBC. Code to reproduce these results can be found at the linked repository (fn 2). This resulted in 1,548 languages classified as “SV”, 331 languages classified as “VS”, and 63 languages classified as “free”.
2.1.2 Nouns and verbs
To measure word length, I counted the number of characters in a word tagged as Noun/Argument or Verb/Predicate [^7] Here I conducted several tests to see if there were different effects for different word types and groupings. “Argument” is a grouping of words tagged as NOUN, PRON, PROPN in the universal tag set, while “predicate” refers to a grouping of VERB and AUX tags. and computed an average length for each class on a per-language basis. [^8] For example, English [ISO 639-3: eng] arguments have an average length of 5.9 characters, while English predicates have an average length of 5.75. English nouns have an average length of 5.9 characters, while English verbs have an average length of 5.8 characters. Irish [ISO 639-3: gle] arguments have an average length of 6.31 characters, while Irish predicates have an average length of 6.29 characters. Irish nouns have an average length of 5.2 characters, while Irish verbs have an average length of 7.9 characters. In the taggedPBC the average number of (unique) arguments per language is 741, and the average number of predicates is 498. The mean length of arguments is 6.71 characters, and the mean length of predicates is 6.82 characters. Conducting a Student’s T-Test showed that this is statistically significant (p < 0.001), indicating that crosslinguistically, arguments are shorter than predicates. All statistics showed a normal distribution.
2.1.3 Relationship between word order and noun/verb lengths
To see whether word order is related to the length of nouns and verbs I conducted a repeated-measures ANOVA with the average lengths of nouns and verbs as the repeated measures and the word order (SV, VS, free) as the between subject factors. Analysis of the results of word order and its relation to length of the different argument and predicate categories is quite striking. First, both SV and VS languages have a highly significant effect of word length, [^9] Additional statistical comparisons and plots are available at the linked repository (Fn 2). whereby arguments are, overall, shorter than predicates crosslinguistically (Figure 3; p < 0.001). [^10] It is also worth noting here that in this corpus words are overall shorter in SV languages than VS languages, which in turn are shorter than those in ‘free’ languages. This displays a general observed trend that languages with free word order tend to have more morphology to identify argument roles and other grammatical relations (Koplenig et al. 2017; Yadav et al. 2020).
Figure 3: Normalized lens of Args, Preds, by word order
Figure 4: Frequency-weighted lens of Ns, Vs, by word order
If we consider only nouns and only verbs, the findings are similar when we normalize the dataset by reducing the vocabulary to unique items. However, if we allow for frequency effects (i.e. we do not reduce the vocabulary to unique items) there is a clear correlation with word order: nouns are longer than verbs in SV languages, but shorter than verbs in VS languages (Figure 4; p < 0.001). This is a major finding that we will return to in the discussion. Setting it aside for now, we turn to two different tests of this measure relating to diachrony: predicting word order change, and the potential interaction with family membership (genealogy/descent) and area (geography).
2.2 Study 2: Predictive validity of word length
Given the observation that relative lengths of nouns and verbs differ in corpora for languages with different word orders, we can ask how predictive this observation is. That is, can we predict the word order of a language based on the relative lengths of its nouns and verbs? If we are strictly interested in identifying basic word order from corpora, it is clear that the N1 ratio already allows us to do so. However, we are also interested in understanding how noun and verb lengths may be related to word order. Being able to predict word order only from noun and verb lengths in a corpus for both a modern language and its ancestral language would indicate that such properties are related to how word order evolves over time.
To investigate this question, for Study 2 I sourced additional corpora of tagged data from two pairs of historical and modern languages in which the modern language changed its word order:
  1. 1.
    Ancient Hebrew (VS) and Modern Hebrew (SV)
  2. 2.
    Classical Arabic (VS) and Egyptian Arabic (SV)
All languages belong to the Semitic branch of the Afroasiatic language family, a necessary constraint given the limited amount of data available for most pre-modern languages. In the case of Hebrew, the UDT has POS-tagged sentences for both Ancient [^11] https://universaldependencies.org/treebanks/hbo_ptnk and Modern [^12] https://universaldependencies.org/treebanks/he_htb varieties. In the case of Classical Arabic I used the Quranic Arabic Corpus [^13] Version 0.4: https://corpus.quran.com and for Egyptian Arabic I used the BOLT Egyptian Arabic Treebank (Maamouri et al. 2018). [^14] https://catalog.ldc.upenn.edu/LDC2018T23 (Discussion Forum) If the lengths of nouns and verbs in these corpora allows us to correctly classify Ancient Hebrew and Classical Arabic as VS and the modern varieties as SV, this would suggest that these are salient properties affecting how word order has developed diachronically. Importantly, given the textual differences between the taggedPBC (New Testament texts) and the test corpora (Biblical, Quranic, modern, and informal texts), such a finding would be significant.
In order to conduct this test, the data for Ancient Hebrew (ISO 639-3: hbo), Modern Hebrew (ISO 639-3: heb), Classical Arabic (cla) [^15] Classical Arabic does not have an assigned ISO 639-3 code or Glottocode. Accordingly, I use the abbreviation cla here, as there is no current language assigned to this code in the ISO 639-3 tables. and Egyptian Arabic (ISO 639-3: arz) were processed and the respective word classes were counted and measured as per the method outlined for Study 1 (each corpus > 400 sentences, see Table 1). The average length of nouns and verbs (including frequency effects) were then used as continuous features for a Gaussian Naive Bayes (GNB) classifier.
Language ISO 639-3 Data Source Sentences Unique nouns Unique verbs
Ancient Hebrew hbo UDT v2.14 409 734 782
Modern Hebrew heb UDT v2.14 882 2068 1297
Classical Arabic cla Quran 6,236 8,243 7,966
Egyptian Arabic arz BOLT 31,688 16,698 10,581
Table 1: Corpora used for testing historical word order change
Three additional features were also used to expand the feature dimensions. The first is a categorical feature: whether a language’s nouns are longer than its verbs. The second is a continuous feature: the difference in length between nouns and verbs. This second measure is related to minimization of complexity, based on the idea that a larger difference between lengths of word classes may effect the likelihood of that particular class being placed first. The third measure is the ratio of noun and verb lengths, and is related to maximization of complexity (via “entropy”), whereby a lower ratio indicates that the particular word classes are less likely to be easily differentiable (and therefore predictable) based on their length.
Using these five features (nounlen, verblen, nlonger, nlen-vlen, nlen/vlen), a GNB classifier was trained on the original combined word order dataset from the taggedPBC containing hand-coded classification of 1,150 languages (Modern Hebrew and Modern Arabic were removed), and then used to predict word order for our pairs of languages: Ancient Hebrew and Modern Hebrew on the one hand, Classical Arabic and Egyptian Arabic on the other. Ancient Hebrew was classified as VS while Modern Hebrew was classified as SV by the trained classifier. Classical Arabic was classified as VS while Egyptian Arabic was classified as SV.
As an additional check, I sourced data from another historical language, Old Irish, as represented in the Codex Paulinus Wirziburgensis, whereby Latin text of St. Paul’s epistles was annotated by Irish scholars during the middle of the eighth century. The dataset (Table 2; Doyle 2018) contains over 3,600 Irish glosses/sentences. As an Indo-European language, Old Irish is widely surmised to have changed its word order to verb-initial from an earlier state, so this check allows us to verify whether the same properties hold in relation to lengths of nouns and verbs. Indeed, the classifier accurately classified Old Irish as VS, based on the data extracted from the Würzburg Glosses.
Language ISO 639-3 Data Source Sentences Unique nouns Unique verbs
Old Irish sga Würzburg 3,647 743 435
Table 2: Irish corpus used for testing historical word order change
Study 2 thus indicates that the relative lengths of nouns and verbs, when frequency is taken into account, can predict word order from corpora unrelated to the type of corpora contained in the taggedPBC, with various implications for our understanding of language processing and language change. This is a striking finding, and I am unaware of any other research that has been able to show predictive validity of a model related to word order change. However, it is possible that this result is conditioned to some degree by the fact that the majority of the test cases are Semitic languages spoken in the same geographical area, so we now turn to an investigation of whether genealogical descent or location provides a better explanation for the correlation of noun and verb lengths with word order.
2.3 Study 3: Testing variance predicted by word length
Dunn et al. (2011) showed that descent from a common ancestor was a better predictor of many grammatical patterns than other explanations. Although Jäger & Wahle (2021) counter this claim with additional data, we can also use the current dataset to see if descent is more explanatory of word order for a given language than the lengths of nouns and verbs. Additionally, in more recent work, Guzmán Naranjo & Becker (2022) show that areal affects are as important as family membership in predicting linguistic phenomena, finding that the results in Dunn et al can be just as easily predicted by language area. This is also a claim that we can test with our dataset. Accordingly, Study 3 can proceed with several (motivated) simplifying observations.
First, for each language we can take the highest level node of family membership that is known based on current scholarship (via Glottolog). [^16] Glottolog (Hammarström et al. 2024; https://www.glottolog.org) contains expert annotations for language family affiliations for over 7,000 languages, along with additional aggregated data from a variety of sources. This is because all modern languages are the same age, and while their pathways/histories may differ, this does not impact whether they can be considered part of the same phylogenetic tree. That is, if descent is more explanatory than word class length, we would expect languages belonging to the same tree to have a higher likelihood of sharing the same word order than not. Second, we can use the macroareas identified by Glottolog as well as the actual GPS coordinates of languages to control for areal affects. Third, we can run various analyses with differing sets of data in order to see how this affects the statistical robustness of the findings. It may be, for example, that using a smaller sample gives a different result than a larger or more diverse one. Here, the more languages we have for a language family or geographical location, the higher the likelihood that any relationship identified is not due to chance. Finally, we can remove languages with “free” word order from consideration, as this leaves us with a binary distinction in word order (SV, VS) that opens up more possibilities for statistical tools.
In particular, a binary dependent variable for word order means we can use hierarchical linear regression (HLR), which allows us to observe the relative impact of particular features that are theoretically relevant to the question at hand by seeing how they influence the performance of successive models. In the case of word order, we can train a regression model with a feature known to be relevant to word order (the N1 ratio), and then subsequent models that incorporate additional features of interest (“language area/location”, “family membership”, “noun/verb lengths”). We can then compare the models to see whether the additional features increase the variance explained in the data via the F-value: an increase in F-value (change) and the significance of this change indicates more/additional variance explained by the particular feature added in a model.
Accordingly, for all languages in the taggedPBC I retrieved the family membership and geographical information from Glottolog, filling in missing data via Ethnologue and other sources. Then I filtered the data on four conditions: a) languages found in Dunn et al that are also present in the taggedPBC (120 languages), b) languages found in the taggedPBC that are in the four families investigated by Dunn et al (662 languages), c) languages in families with more than 75 members in the taggedPBC (1182 languages), and d) languages in families with 2 or more members in the taggedPBC (1,805 languages). For each of these conditions, I ran an HLR, adding features of interest progressively to see how they impacted the fit of the successive models and whether particular features account for additional variance.
The base model (Model 1) uses the N1 ratio as the primary feature for classification. This is because we know from comparison with expert determinations of word order (see Ring 2025) that the N1 ratio significantly correlates with the classification of languages as SV or VS. Due to the strong relationship observed between language area and typological properties of language (see Guzmán Naranjo & Becker 2022), Model 2 then includes GPS coordinates and macroarea identifiers for each language as features. Model 3 additionally incorporates descent (language family membership) as a feature (as per Dunn et al. 2011), and Model 4 includes the average lengths of nouns and verbs (as per the relationship identified in this paper). If descent is more explanatory than the lengths of nouns/verbs, this should be clear from the differences between the respective models for each condition. Results are displayed in Table 3.
a) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 120 45.27 6.4e-10
2 (Lg area) 120 24.29 1.3e-14 12.78 2.87e-07 **
3 (Family) 120 20.02 2.7e-14 2.05 0.154
4 (N/V len) 120 15.75 2.9e-14 3.17 0.045 *
b) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 662 189.99 3.6e-38
2 (Lg area) 662 95.65 4.1e-64 50.07 3.7e-29 **
3 (Family) 662 78.23 2.4e-64 5.79 0.016 *
4 (N/V len) 662 65.43 2.6e-71 21.30 1.0e-09 **
c) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 1182 378.83 2.1e-73
2 (Lg area) 1182 210.38 3.3e-136 116.99 2.6e-66 **
3 (Family) 1182 169.09 1.3e-135 2.72 0.099
4 (N/V len) 1182 127.47 2.3e-139 14.03 9.4e-07 **
d) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 1805 535.41 6.1e-104
2 (Lg area) 1805 178.89 3.8e-129 46.52 5.8e-29 **
3 (Family) 1805 143.49 2.3e-128 1.66 0.197
4 (N/V len) 1805 108.16 2.2e-132 14.46 5.8e-07 **
Table 3: Hierarchical Linear Regressions under 4 conditions
From these analyses we can see that, as expected, the N1 ratio significantly predicts word order under all conditions. Additionally, language area is a significant predictor of word order that adds to the strength of the classification. Language family (descent) does not add additional predictive power to our model under most conditions (over language area), whereas the length of nouns and verbs does. This analysis shows that the length of nouns and verbs in a language accounts for more variance and at the same time is more significant than that language’s family membership in determining its word order classification, under all test conditions. When we remove language area from the linear regression this is still the case, [^17] See additional statistical results at the linked repositories (Fn 2; Fn 5). indicating that while language area is a stronger predictor than family membership, lengths of nouns and verbs is yet a stronger one.
This largely aligns with Jäger and Wahl’s (2021) finding regarding the (lower) correlation between descent and linguistic structures in comparison to other explanations, with the caveat that Jäger and Wahl included language isolates, which is likely to greatly diminish the affect of descent. If a large number of languages in a sample are not related to others, clearly descent can have only a minimal impact on word order. By including only language families with two or more members in the taggedPBC (which excludes 111 languages), I show that descent does not explain much variance above and beyond language location/area, in contrast with noun and verb lengths. This also suggests that language family/descent is highly correlated with linguistic area, which conforms with expectations related to human history and migration. Additionally, the amount of variance explained seems to depend on the number of languages (and language families) in the sample, highlighting the need for more data for languages from under-represented language families.
3 Discussion
That arguments overall are shorter in languages of the taggedPBC, across all word orders, suggests that there are certain (non-Noun) arguments affecting these observations. Particularly, the effect of PRON seems quite likely, whereby a tendency for pronouns to be shorter crosslinguistically, along with a preference for pronouns as arguments, may impact the observation that arguments are shorter overall than predicates. This pattern aligns with the processing claim that “heavier” words (or word classes in this case) occur later in a sentence, partly validating Hypothesis 1 of study 1. The counter-argument, also from the data, is that nouns are longer than verbs in SV languages and shorter than verbs in VS languages when frequency is considered, validating Hypothesis 2 of study 1. This suggests that speakers prefer to place heavier (more “contentful”, or perhaps “novel”) constituents before lighter ones crosslinguistically in many cases.
These seemingly antithetical results partially validate both the “efficiency”-based theory and the “surprisal” theory, highlighting the interaction of two pressures: one whereby speakers prioritize lighter constituents (possibly for the purposes of production/processing efficiency), and one where they prioritize heavier constituents (possibly for informational purposes). The two perspectives may in fact be looking at the same data from different angles. On the one hand, grouping all arguments into a single class highlights the overall impact of shorter strings, since the most frequently used arguments (pronouns, demonstratives) are typically shorter crosslinguistically. On the other, more “content-ful” word types that aid planning and inference or predictability (nouns, verbs) will typically be longer crosslinguistically.
These two pressures can be (and, in fact, have been) considered as two different kinds of “efficiency”: production efficiency vs communicative efficiency. Under this view, production efficiency is related to length of words (as per Wasow 2002, 2022; Hawkins 2014) - speakers use shorter words whenever they can to efficiently produce utterances (see also Mahowald et al. 2013). The counter-pressure, communicative efficiency, is related to making a speaker’s utterances more readily accessible/interpretable by recipients (as per Levshina & Moran 2021; Levshina 2023) - speakers prioritize words that contain more meaningful, informative content in order to aid communication. Once context has been established, these longer words can be replaced by more efficient (shorter) terms such as demonstratives or pronouns. [^18] For example, there is a common crosslinguistic pattern of using a noun first to establish reference, then switching to pronouns, i.e. "that tree, it is big, it will fall…" See also discussion of the (lack of) encoding of subjects crosslinguistically in Berdicevskis et al. (2020).
I would argue, however, that conceptualizing “efficiency” in these two distinct ways, as currently done in the literature, runs the risk of confusing the two. It is in fact not efficient (from a processing perspective) to maximize information, as this also tends to maximize complexity. Instead, I propose a different framing of these principles.
3.1 A “Min-Max” theory of processing
The current proposal re-formulates or refines Hawkins’ PGCH to incorporate (increased) complexity as a motivation for the order of constituents, bearing some resemblance to proposals in Levshina (2023). I describe this as (to borrow a term from computer gaming) [^19] A “min-maxer” in gaming terminology is an individual who minimizes their effort while maximizing their gains. This can be in terms of focusing their efforts on activities in a game that give the best returns, or by minimizing survival stats on a character in order to maximize damage (as in a “glass cannon” build for ranged attackers in an rpg). the “Min-Max” theory of processing (schematized in Figure 5), whereby speakers seek to minimize their processing effort while maximizing the information in their utterances. These are competing/parallel pressures that speakers are constantly balancing in production (or “performance”) in order to choose the best strategy for their immediate goals. As speakers produce variants along a cline of grammaticality, these variants compete on the basis of multiple factors, including how well such variants support Min-Max-ing in relation to other variants, leading to language change.
Figure 5: Min-Max schema
These two competing pressures offer a more complete (and parsimonious) explanation for word order change (and possibly other kinds of linguistic change) than existing theories of language. In one possible scenario the frequent use of short pronominal forms encourages drift toward an argument-initial order, and nouns subsequently lengthen as they provide more information to speakers and move toward an initial position. This, combined with the preponderance of shorter pronominal forms crosslinguistically, may provide an explanation for why the majority of languages reported in the world are SV. Conversely, if a language has or develops longer verbs, the preference for these heavier constituents to occur first could put pressure on the language to maintain or develop predicate-initial order, which would explain why some VS languages are quite stable. In each case the direction of development would depend on a number of factors that have been implicated in discussions of language change and grammaticalization (see Heine & Kuteva 2002; Matras 2011; Narrog & van der Auwera 2011). However, the effect of word length is clearly a major part of the explanation for why we can observe (and predict) such a difference in word order between Ancient and Modern Hebrew as well as Classical and Egyptian Arabic via their respective corpora.
The existence of multiple competing pressures affecting the development of language, as described by this Min-Max theory, is supported by work in clinical psychology (e.g. Rezaii et al. 2023) and language modeling (e.g. Hahn & Xu 2022), as well as findings from information theory (Zaslavsky 2020; Tucker et al. 2025). In a study of patients with nonfluent aphasia, Rezaii et al. (2023) showed that patients with this condition compensate for their deficit in producing complex grammatical structures by increasing the information content of their utterances, dropping high-frequency (function) words in favor of low-frequency (content) words (i.e. “maximizing information”). By modeling variations in grammatical structure based on corpora from 80 languages, Hahn & Xu (2022) found that the resulting models were less optimized for dependency length (DL) and information locality (IL) than the original structures of each language, suggesting that languages evolve to efficiently balance the two considerations of DL and IL. [^20] Notably, Hahn & Xu propose a theory that the interaction between Dependency Length (DL) and Information Locality (IL) accounts for word order variation and change, but neither of the two concepts, as expressed, make claims about directionality/order, only about the relationships between (internal) constituents. In contrast, the Min-Max theory makes the claim that speakers balance information content (as reflected by longer nouns in SV languages and longer verbs in VS languages) with processing/production concerns, and that this interaction is what drives the evolution of word order. As nouns lengthen or verbs shorten in a language, this will affect their placement in a stream, which then impacts the overall linear order observed in surface structures. That is, the Min-Max theory makes a universal claim regarding language-external pressures, while DL and IL are language-internal concepts. These and other studies indicate that language users are sensitive to the pressures of minimizing processing and maximizing information content.
The sensitivity of language users to processing efficiency and information content has also been a major topic of investigation in information theory, which is concerned in part with the encoding and decoding of information streams. One information-theoretic proposal is that semantic systems evolve following an “Information Bottleneck” principle (IB; Tishby et al. 1999), whereby “optimal representations satisfy a tradeoff between compressing X, i.e. minimizing the representational complexity, and maximizing the relevant information about Y” (Zaslavsky 2020: 14). Zaslavsky shows that IB can account for the development of color naming systems in language, and additional research shows the utility of IB in machine learning and distributional models of semantics (Slonim & Tishby 2000).
Implications of this interaction have most recently been examined in agent-based emergent learning simulations with the addition of a “utility” principle (Tucker et al. 2025) – the interaction between “utility”, “informativeness”, and “complexity” allows more human-like communication to emerge among agent models in their study. Such a finding is very much supported by the results reported here, though I would argue that the Min-Max theory offers a more parsimonious account of human behavior as related to language specifically. While the current corpus evidence applies at the word level in relation to basic word order, it seems likely that the minimization of processing difficulty would interact with the maximization of information content at multiple (hierarchical) levels. Speakers min-max in order to select words in a stream (“informativeness” and “complexity” as per Tucker et al.), while simultaneously min-maxing to select relevant structures related to their communicative goals (Tucker et al.’s “utility”).
As noted above, the Min-Max theory also bears some resemblance to Levshina’s (2023) proposals, described in terms of communicative efficiency and accessibility. However, such indicators cannot fully explain certain patterns in grammar such as “harmonic” word order. A recent proposed explanation for such patterns involves frequency of word classes, whereby Mansfield & Krapp (2025) suggest that “harmonic ordering is really a frequency effect: as replication converges on a phrase structure with a consistent linear order, the most frequent word class tends to be at one edge.” Although they did not test token length in their model, it is notable that in Study 1 (§2.1 above) frequency is a major factor in identifying the relationship between length and word order, suggesting that the most frequent word class in their model may also be the longest element of the domain they are testing, which would be driven by the maximization of information principle. The Min-Max theory might also provide an explanation for non-parallel (hierarchical) harmonic word order, which is not explained by their model.
The Min-Max theory may additionally explain why word forms remain long (or lengthen), as described for particular languages (see Cristofaro 2021; Haspelmath 2021; Petré 2017). Here, again, the maximization of information principle would support the development or retention of particular (long) word forms, as speakers balance this need against the minimization of processing difficulty. As noted above, this proposal is supported by evidence from color-naming systems, which show patterns in keeping with IB (Zaslavsky 2020). The Min-Max theory would suggest that this principle does not only hold for “forms” (of, i.e. color words), but also governs “behavior”: speakers behave in ways that are consistent with maximizing information in streams (via word order), while minimizing processing effort.
Such a theory can be conceived of as a universal (underlying) pressure that applies generally to how language is realized. From an innateness perspective, speaker choices impacted by such pressures would lead to a re-analysis or re-formation of rules that guide surface realizations. From a usage-based perspective such pressures would lead to the development of new surface variants. These are not necessarily opposing views (see Rastelli 2025), and the outcomes are quite similarly dependent on language being viewed as a complex adaptive system, whereby diachronic changes may have a combination of both “result-oriented” and “source-oriented” explanations (see Cristofaro 2019; Schmidtke-Bode & Grossman 2019).
Another observation from the results reported here is that information content seems to trump length. We find that long nouns are used with more frequency in SV languages, while long verbs are used with more frequency in VS languages. This indicates that (maximizing) information content is more important to speakers than (minimizing) processing complexity. This could explain the “agent/subject-first preference” in experiments (Culbertson et al. 2012; Sauppe et al. 2023), as the “information maximization” pressure wins out in cases where this most efficiently meets communicative goals.
Further, the data shows that all languages have a relatively high proportion of sentences in which the argument occurs before the predicate. In VS languages, the proportion of argument-initial to predicate-initial sentences is 1/2 or more (min: 0.56 n/v, 0.65 a/p), highlighting that even VS languages still place the noun/argument first in many constructions. Similarly, while all SV languages have a high proportion of N1 sentences (max 2.43 n/v, 7.17 a/p), the number of V1 constructions is non-zero, supporting a gradient approach to word order (Levshina et al. 2023).
The fact that the length of particular word classes can predict the basic word order of a language is a novel finding, leading to some general predictions: if, on average, nouns are longer than verbs in a language, it is highly likely to have SV word order. If, on the other hand, verbs are longer, it is highly likely to have VS word order. The effect of frequency cannot be ignored either, highlighting the importance of conducting large-scale corpus based studies to investigate this question. Since languages are used by speakers and such speakers use certain words more than others, we can only observe this frequency effect (or “weight”) of word class length via corpora.
The idea that speakers balance multiple considerations in production is not new (Pike 1967; Jackendoff & Pinker 2005; Seržant & Moroz 2022; Hörberg & Sjons 2023) but what is novel is the idea that all humans experience a specific universal pressure to balance production/processing efficiency with information content. This pressure is observable crosslinguistically, and would apply at multiple processing levels, regardless of whether a speaker is accessing a universal grammar (“innate”) or inferring structure on the basis of the input they receive (“usage”). Following on from Zipf’s and subsequent observations, there is no doubt that the length of strings is an important factor in processing streams, but clearly not all strings are created or treated equally by speakers. The studies presented here show that nouns are treated differently (or at least have a different distributional relationship) than other kinds of arguments, which leads to different realizations of word order crosslinguistically. The pressures identified in this paper have wide-ranging implications and applications for any conceptualization of how language operates, whether in terms of understanding the intricacies of individual linguistic scenarios, designing intelligent systems that use language, or developing clinical interventions to assist speakers in various ways.
3.2 Limitations and future directions
One limitation in this set of studies is the use of automated processes to identify part of speech tags in the data. Ideally, all the data should be hand-annotated for parts of speech, which would allow for better comparison of the different word categories across languages. Given the time consuming nature of this task and the need for expert annotators, this is not yet possible, but ongoing research could address this by sourcing experts to annotate the data. Although I have shown (in Ring 2025 and via the linked repository) that the translation/alignment method does correspond well to hand-tagged data, it may be that this could be improved by tagging a small amount of data for each language in the PBC in order to train POS taggers that perform better than the current translation method.
Another limitation is that the dataset used, while exceedingly large in comparison to previous efforts, still only represents about a quarter of the world’s languages. Since the majority of the world’s languages are not represented here, it is not clear that these findings can be fully generalized to all languages of the world, though I have shown in section 2.3 that much can be learned from sampling. Additionally, the contrast between intransitive SV/VS/free word order is somewhat of an oversimplification of word order patterns. Given the current obstacles to annotation, it is a necessary simplification for this study, however future work could expand to a more diverse set of word/constituent orders and potential relationships with word length and/or complexity. Given the nature of the dataset used, it is also possible to pursue a gradient approach to word order, as advocated by Levshina et al. (2023). The issue of word complexity could additionally use further expansion to cover a broader range of possibilities than simply the length of a word based on a set of (roman) characters.
A final note is that our conclusions are only as good as the data that we have available to us. The taggedPBC is a significant step in the right direction, toward a representative sample of large tagged corpora for languages of the world. Systematic investigation of this dataset can give us insight into language typology (the current state of language), as well as pressures that impact this realization in diachrony, both in terms of how languages have developed (their history) and how they might develop (their future). On the basis of this investigation I have proposed a parsimonious “Min-Max” theory of processing and information structure that underlies language evolution, whereby speakers minimize processing effort while maximizing information gain. This proposal is in keeping with at least one existing information-theoretic proposal (IB) as well as theories of processing (efficiency; PGCH; surprisal) and can explain the correlation between the length of nouns/verbs and basic word order, given that these features allow us to predict the word order of a language. It is as yet unclear whether such a principle would apply at multiple hierarchical levels within language, which is an area that deserves further careful investigation with datasets such as the taggedPBC. For now, I have demonstrated that noun/verb lengths are more explanatory than descent or geography and that for languages with corpus data, word length predicts word order.